# Training a Foundation Model for Materials on a Budget

### v1

@@ Abstract
Foundation models for materials modeling are advancing quickly, but their training remains expensive, often placing state-of-the-art methods out of reach for many research groups. We introduce Nequix, a compact E(3)-equivariant potential that pairs a simplified NequIP design with modern training practices, including equivariant root-mean-square layer normalization and the Muon optimizer, to retain accuracy while substantially reducing compute requirements. Built in JAX, Nequix has 700K parameters and was trained in 500 A100-GPU hours. On the Matbench-Discovery and MDR Phonon benchmarks, Nequix ranks third overall while requiring less than one quarter of the training cost of most other methods, and it delivers an order-of-magnitude faster inference speed than the current top-ranked model. We release model weights and fully reproducible codebase at https://github.com/atomicarchitects/nequix.
@@ 1 Introduction
We pursue an orthogonal goal to scaling: a lower computational cost recipe that preserves strong downstream accuracy. Concretely, we revisit a simplified E(3)-equivariant architecture based on NequIP (Batzner et al., 2022) with modern training practices: root-mean-square layer normalization for stability, dynamic batching to maximize GPU utilization, and optimizer choices inspired by “speedrunning” deep learning workflows (Jordan et al., 2024a). The resulting model, Nequix, has 700K parameters and can be trained in 500 GPU hours, while remaining competitive with larger and more costly to train models on Matbench-Discovery and other phonon prediction tasks.
@@ 2 Methods
Figure 1: (a) Nequix architecture, a simplified version of NequIP (Batzner et al., 2022), with a species-independent residual connection and layer normalization. (b) Combined performance scores of compliant models on the Matbench-Discovery (unique prototypes subset). (c) Available published training times of current compliant models.
@@
Nequix is implemented in JAX (Bradbury et al., 2018) with Equinox (Kidger and Garcia, 2021), taking advantage of just-in-time compilation and efficient automatic differentiation. Following standard energy-conserving MLIP practice (Fu et al., 2025), forces are obtained as the negative energy gradient with respect to atomic positions, $\nabla_{\mathbf{r}}E$, and stresses as the energy derivative with respect to strain, normalized by volume, $\sigma=V^{-1}\,\partial E/\partial\varepsilon$, where $E$ is the predicted total energy of the system, $\mathbf{r}$ is an atom position, $\varepsilon$ is the strain tensor, $V$ is the simulation cell volume, and $\sigma$ denotes the stress tensor.
@@
Materials radius graphs vary widely in their numbers of nodes and edges due to differences in unit-cell size and atom count. With fixed-size batching, the largest graphs dictate memory usage, leaving GPUs underutilized for most batches. To keep batch workloads more uniform while respecting memory limits, we use dynamic batching (Speckhard et al., 2025): each batch is filled up to caps on the total nodes and edges. We set these caps to $1.1\times$ (ideal batch size) $\times$ (dataset-average nodes-per-graph and edges-per-graph), respectively. Batches are then padded to these caps to fulfill the static shape requirements of JAX (Godwin* et al., 2020).
@@
The final Nequix model is trained for 100 epochs on MPtrj (Jain et al., 2013; Deng et al., 2023), of which we hold out 5% for validation. More details on the training settings are provided in Sec. A.1. The model was trained on 4 NVIDIA A100 GPUs in 125 hours, for a total cost of 500 GPU hours.
@@ 4 Conclusion
We presented Nequix, an E(3)-equivariant interatomic potential that pairs a simplified NequIP architecture with modern training practices. Our results show that Nequix achieves competitive accuracy on Matbench-Discovery and the MDR phonon benchmark at less than one quarter of the reported training cost of many contemporaries. This resource-efficient recipe provides a practical alternative to large-scale foundation models and helps broaden access to high-quality atomistic modeling in settings with more limited compute. We release trained weights and a JAX codebase to streamline reuse and extension.
@@
We thank Mit Kotak for helpful discussions and insights. This work was supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) and by DOE ICDI grant DE-SC0022215. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ERCAP0033254.

### v2

@@ Abstract
Foundation models for materials modeling are advancing quickly, but their training remains expensive, often placing state-of-the-art methods out of reach for many research groups. We introduce Nequix, a compact E(3)-equivariant potential that pairs a simplified NequIP design with modern training practices, including equivariant root-mean-square layer normalization and the Muon optimizer, to retain accuracy while substantially reducing compute requirements. Nequix has 700K parameters and was trained in 100 A100 GPU-hours. On the Matbench-Discovery and MDR Phonon benchmarks, Nequix ranks third overall while requiring a 20 times lower training cost than most other methods, and it delivers two orders of magnitude faster inference speed than the current top-ranked model. We release model weights and fully reproducible codebase at https://github.com/atomicarchitects/nequix.
@@ 1 Introduction
We pursue an orthogonal goal to scaling: a lower computational cost recipe that preserves strong downstream accuracy. Concretely, we revisit a simplified E(3)-equivariant architecture based on NequIP (Batzner et al., 2022) with modern training practices: root-mean-square layer normalization for stability, latest custom CUDA kernels (Bharadwaj et al., 2025), and optimizer choices inspired by “speedrunning” deep learning workflows (Jordan et al., 2024a). The resulting model, Nequix, has 700K parameters and can be trained in 100 GPU hours, while remaining competitive with larger and more costly to train models on Matbench-Discovery and other phonon prediction tasks.
@@ 2 Methods
Figure 1: (a) Nequix architecture, a simplified version of NequIP (Batzner et al., 2022), with a species-independent residual connection and layer normalization. (b) Combined performance scores of compliant models on the Matbench-Discovery (unique prototypes subset), collected on 2025-08-17. (c) Available published training times of current compliant models.
@@
Nequix is implemented in both JAX (Bradbury et al., 2018; Kidger and Garcia, 2021) and PyTorch(Paszke et al., 2019; Ansel et al., 2024), taking advantage of just-in-time compilation and efficient automatic differentiation. Following standard energy-conserving MLIP practice (Fu et al., 2025), forces are obtained as the negative energy gradient with respect to atomic positions, $\nabla_{\mathbf{r}}E$, and stresses as the energy derivative with respect to strain, normalized by volume, $\sigma=V^{-1}\,\partial E/\partial\varepsilon$, where $E$ is the predicted total energy of the system, $\mathbf{r}$ is an atom position, $\varepsilon$ is the strain tensor, $V$ is the simulation cell volume, and $\sigma$ denotes the stress tensor.
@@
Materials radius graphs vary widely in their numbers of nodes and edges due to differences in unit-cell size and atom count. With fixed-size batching, the largest graphs dictate memory usage, leaving GPUs underutilized for most batches. To keep batch workloads more uniform while respecting memory limits, we use dynamic batching (Speckhard et al., 2025): each batch is filled up to caps on the total nodes and edges. We set these caps to $1.1\times$ (ideal batch size) $\times$ (dataset-average nodes-per-graph and edges-per-graph), respectively. Batches are then padded to these caps to fulfill the static shape requirements of JAX (Godwin* et al., 2020). This capability is only used for the JAX implementation.
@@
GPU kernels
There has been recent work (Bharadwaj et al., 2025; NVIDIA, 2024; Tan et al., 2025; Lee et al., 2025) on writing custom GPU kernels for the expensive equivariant tensor product operation (Xie et al., 2025). These kernels fuse the tensor product and the outer gather-scatter from the message passing step into a single GPU kernel. This avoids storing costly edge-based intermediates in GPU memory, improving both runtime and memory usage.
@@
The final Nequix model is trained for 100 epochs on MPtrj (Jain et al., 2013; Deng et al., 2023), of which we hold out 5% for validation. More details on the training settings are provided in Sec. A.1. The model was trained on 2 NVIDIA A100 80 GB GPUs in 50 hours, for a total cost of 100 GPU hours.
@@ 3.3 Inference speed
Finally, we compare inference speed with existing interatomic potentials by using ASE Larsen et al. (2017) to run calculations on diamond across varying unit cell sizes, timing the calculations for each model. [^1] See https://github.com/mitkotak/matbench-speed for inference speed benchmarking code.. We run each model in the default configuration in which it used within its ASE calculator, with compilation and kernels wherever specified in the documentation. See Section A.2 for more information on benchmarking setup.
@@ 4 Conclusion
We presented Nequix, an E(3)-equivariant interatomic potential that pairs a simplified NequIP architecture with modern training practices. Our results show that Nequix achieves competitive accuracy on Matbench-Discovery and the MDR phonon benchmark at less than one quarter of the reported training cost of many contemporaries. This resource-efficient recipe provides a practical alternative to large-scale foundation models and helps broaden access to high-quality atomistic modeling in settings with more limited compute. We release trained weights and a JAX/PyTorch codebase to streamline reuse and extension.
@@
This work was supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) NSF Graduate Research Fellowship program under Grant No. DGE-1745302, and by DOE ICDI grant DE-SC0022215. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ERCAP0033254.
@@ A.1 Training and model configuration
A.2 Inference Configuration
We summarize all of the inference configurations in Table A.2. We use MPTrj versions for all of the models except NequIP where the MPtrj model did not have kernel support, so we instead use the OMat24 model after confirming that they have the same hyperparams. For compatibility with eSEN, we use the older fairchem OCP calculator, and expect the latest version to have less overhead.
Table A.2: Inference configuration for the ASE benchmarking setup
Model Compile Kernels
eSEN-30M $\times$ $\times$
MACE-MP0 torch.compile cuEquivariance
Nequix torch.compile openEquivariance